# ChartGalaxy: A Dataset for Infographic Chart Understanding and Generation

### v3

@@ Appendix A Chart-rich Websites
Table 5: Chart-rich websites and their terms of service.
Name URL Terms of Service
Pinterest pinterest.com Link [^1] https://policy.pinterest.com/en/terms-of-service
Statista statista.com Link [^2] https://www.statista.com/getting-started/publishing-statista-content
Visual Capitalist visualcapitalist.com Link [^3] https://licensing.visualcapitalist.com/recent-changes-to-visual-capitalist-licensing/
World Statistics world-statistics.org Link [^4] https://world-statistics.org/terms-of-use.php
Chart Porn chartporn.org Link [^5] https://chartporn.org/about/
Cool Infographics coolinfographics.com/blog Link [^6] https://coolinfographics.com/about
Daily Infographics dailyinfographics.eu Link [^7] https://dailyinfographic.com/terms
Data Pointed datapointed.net Link [^8] http://www.datapointed.net/about/
DataRep datarep.com Link [^9] https://www.datarep.com/website-terms-of-use/
Flickr flickr.com Link [^10] https://www.flickr.com/help/terms
FlowingData flowingdata.com Link [^11] https://flowingdata.com/about/
Infographics Archive infographicsarchive.com Link [^12] https://www.infographicsarchive.com/about/
Information is Beautiful informationisbeautiful.net Link [^13] https://informationisbeautiful.net/about
National Post nationalpost.com Link [^14] https://www.postmedia.com/terms-and-conditions
Pikto Chart piktochart.com Link [^15] https://piktochart.com/terms-of-use/
SCMP scmp.com Link [^16] https://www.scmp.com/terms-conditions
Visme visme.co Link [^17] https://www.visme.co/about/
Visual Cinnamon visualcinnamon.com Link [^18] https://www.visualcinnamon.com/about/
Chit Chart chitchart.com Link [^19] https://chitchart.com/terms-use
Appendix B Chart Types and Variations
We provide a full list of 75 chart types and 330 chart variations.
B.1 Chart Types
@@
We summarize chart types observed in the collected real infographic charts, drawing on two existing taxonomies: Data Viz Project [45] and Datylon [46]. The diversity of chart types ensures that our synthetic infographic charts can adapt to a wider range of scenarios and data representations, making them valuable for model training and evaluation. The full list of 75 chart types is in Figs. 5-20.
B.2 Chart Variations
For each chart type, we extract chart variations from real-world infographic charts, featuring diverse visual styles, e.g., element shapes, icon placement, hand-drawn styles, and 3D styles. Therefore, we provide illustrations of the variations under each chart type (Figs. 5-20).
B.3 Mapping Rules
@@
We determine chart types by analyzing the data attributes and their characteristics. First, we apply rule-based mapping (Table LABEL:tab:chart_mapping_rules) that identifies candidate chart types based on data attribute combinations. If multiple candidates remain, we instruct Gemini-2.0-Flash to select the best fit by evaluating the compatibility between the data and these chart types. The selection prompt, shown below, includes candidate chart types, their descriptions, and key data statistics to guide the decision.
@@
Figure 5: 75 chart types and 330 chart variations (Part 1).
Figure 6: 75 chart types and 330 chart variations (Part 2).
Figure 7: 75 chart types and 330 chart variations (Part 3).
Figure 8: 75 chart types and 330 chart variations (Part 4).
Figure 9: 75 chart types and 330 chart variations (Part 5).
Figure 10: 75 chart types and 330 chart variations (Part 6).
Figure 11: 75 chart types and 330 chart variations (Part 7).
Figure 12: 75 chart types and 330 chart variations (Part 8).
Figure 13: 75 chart types and 330 chart variations (Part 9).
Figure 14: 75 chart types and 330 chart variations (Part 10).
Figure 15: 75 chart types and 330 chart variations (Part 11).
Figure 16: 75 chart types and 330 chart variations (Part 12).
Figure 17: 75 chart types and 330 chart variations (Part 13).
Figure 18: 75 chart types and 330 chart variations (Part 14).
Figure 19: 75 chart types and 330 chart variations (Part 15).
Figure 20: 75 chart types and 330 chart variations (Part 16).
Appendix C Layout Templates and Examples
C.1 Layout Templates
We summarize 68 layout templates from real infographic charts. The full list of these templates is in Figs. 21 and 22.
C.2 Examples
@@
Appendix D Extended Evaluation and Results
D.1 Instruction Dataset for Infographic Chart Understanding
@@
We fine-tune two LVLMs in our experiments. The first is InternVL3-8B [68], which combines the Qwen2.5-7B language model with the InternViT-300M-448px-V2.5 visual encoder. The second is Qwen2.5-VL-7B [69], which integrates the same Qwen2.5-7B language model with a Vision Transformer architecture optimized by Qwen. These models are fine-tuned on our instruction dataset for 2 epochs using a global batch size of 128. For parameter-efficient tuning, we utilize LoRA [84] with a rank of 8. This training process is conducted on 8 NVIDIA 4090D 48G GPUs, leveraging Deepspeed ZerO for parallelized training. For both InternVL3-8B and Qwen2.5-VL-7B, the learning rate is set to $5\times 10^{-5}$, and we employ a cosine learning rate scheduler with a warm-up ratio of $0.1$. The LoRA parameters, $\alpha$ and $r$, are both set to 32.
@@
D.2 Benchmarking Infographic Chart Code Generation
@@
Table 8 lists the 17 LVLMs along with their API names used in this experiment. Since our task involves generating executable code with relatively long outputs, we conducted a small-scale pilot study to assess the basic code generation capability of each model. Based on this, we excluded models that consistently failed to produce meaningful or complete outputs under our task setting—for example, Phi-4—due to their limited capacity or inability to handle long sequences. We use greedy decoding (temperature $\tau=0$) across all models to ensure deterministic outputs. To maximize the chance of obtaining complete and executable code, we configure each model to generate as many tokens as possible, setting the maximum generation length to $min(16384,A)$, where $A$ denotes the model’s maximum generation limit. This helps mitigate the risk of incomplete outputs, which result in non-executable code.
@@
Gemini-2.5-Pro [73] Proprietary gemini-2.5-pro-preview-05-06
Gemini-2.5-Flash [77] Proprietary gemini-2.5-flash-preview-04-17
Claude-3.7-Sonnet [75] Proprietary claude-3-7-sonnet-20250219
GPT-4.1 [74] Proprietary gpt-4.1
GPT-4.1-mini [74] Proprietary gpt-4.1-mini
GPT-4.1-nano [74] Proprietary gpt-4.1-nano
OpenAI-o4-mini [76] Proprietary o4-mini
OpenAI-o3 [76] Proprietary o3
OpenAI-o1 [78] Proprietary o1
GPT-4o [79] Proprietary gpt-4o-2024-11-20
Doubao-1.5-Vision-Pro [80] Proprietary Doubao-1.5-vision-pro-32k
Moonshot-v1-Vision [81] Proprietary moonshot-v1-32k-vision-preview
Llama-4-Maverick-17B [82] Open-Source chutesai/Llama-4-Maverick-17B-128E-Instruct-FP8
Llama-4-Scout-17B [82] Open-Source chutesai/Llama-4-Scout-17B-16E-Instruct
Qwen2.5-VL-72B [69] Open-Source Qwen/Qwen2.5-VL-72B-Instruct
Qwen2.5-VL-32B [69] Open-Source Qwen/Qwen2.5-VL-32B-Instruct
InternVL3-78B [68] Open-Source internvl3-78b
@@
For the high-level score, we employ GPT-4o [79] to assess the visual similarity between the PNG image rendered by the generated code and the ground-truth one. We instruct GPT-4o to evaluate the similarity along six dimensions: data element, layout, text, image, color, and validity. The model outputs a score for each of the six dimensions, which are then summed to produce a total score ranging from 0 to 100. The detailed prompt for this evaluation is provided in Supp. E.2.
The low-level score evaluates the fine-grained similarity between SVG elements of the rendered chart and the corresponding ground-truth chart. This evaluation is conducted through three steps: 1) decomposing both charts into SVG elements, 2) matching elements between the two charts, and 3) computing similarity metrics based on the matching results [70, 72]. Algorithm 1 presents the pseudo-code for the matching procedure. The full implementation details are available in our publicly accessible code repository [^21] https://github.com/ChartGalaxy/ChartGalaxy.
@@
1:$gt\_leafs$: Ground truth SVG elements.
2:$pr\_leafs$: Predicted SVG elements.
3:$gt\_matched$: Array for ground truth matches (init with -1).
4:$pr\_matched$: Array for prediction matches (init with -1).
@@
6:$m\leftarrow|gt\_leafs|$, $n\leftarrow|pr\_leafs|$
7:$CostMatrix\leftarrow\text{ZeroMatrix}(m,n)$
@@
13:$(rows,cols)\leftarrow\text{HungarianAlgorithm}(CostMatrix)$triangleright Returns optimal row-column pairs
@@
15:     if $CostMatrix[i][j]\leq 1$ AND $gt\_matched[i]=-1$ AND $pr\_matched[j]=-1$ then
@@
The text and image metrics evaluate the similarity of generated text and image elements, respectively, by averaging the similarity scores over all matched pairs of text and image elements. Unmatched text and image elements in ground-truth charts are assigned a similarity score of 0 to penalize generation failures. For text elements, similarity is computed using the character-level Sørensen-Dice coefficient, defined as twice the number of overlapping characters divided by the total number of characters in the two strings [70]. For image elements, similarity is measured using the CLIP embedding-based similarity [44].
The color, position, and size metrics assess visual consistency across matched elements with respect to their respective attributes. The color metric employs the CIEDE2000 formula [85] to measure the perceptual difference between the colors of matched elements. The position and size metrics are defined as follows:
@@
Gemini-2.5-Pro [73] 6,662.45 1,674.01 85.21
GPT-4.1 [74] 5,237.78 1,675.36 80.00
Claude-3.7-Sonnet [75] 5,870.50 1,552.48 79.91
GPT-4.1-mini [74] 7,656.71 2,488.46 79.69
OpenAI-o4-mini [76] 2,886.75 735.19 75.97
Gemini-2.5-Flash [77] 6,508.22 1,934.79 75.55
OpenAI-o1 [78] 2,662.11 1,018.11 74.69
OpenAI-o3 [76] 3,114.74 999.66 74.22
GPT-4o [79] 2,962.72 613.42 65.67
GPT-4.1-nano [74] 2,387.05 1,306.72 60.06
Doubao-1.5-Vision-Pro [80] 3,878.70 1,347.10 47.11
Moonshot-v1-Vision [81] 3,087.17 791.16 44.39
@@
Llama-4-Maverick-17B [82] 3,460.76 856.62 61.29
Qwen2.5-VL-72B [69] 3,512.20 1,236.29 57.09
InternVL3-78B [68] 3,376.52 758.71 55.07
Llama-4-Scout-17B [82] 3,247.25 826.56 51.91
Qwen2.5-VL-32B [69] 4,960.82 1,264.43 46.48
@@
Prompting methods. In addition to the direct prompting method, we evaluate three alternative prompting strategies following prior work [40, 70]:
@@
  • SelfReflection, where the model is provided the ground-truth chart, previously generated code from direct prompting, and the rendered chart, and is instructed to revise the given code.
@@
Llama-4-Maverick-17B [82] Direct 100.00 78.24 58.51 59.53 66.35 74.00 47.74 64.06 58.42 61.24
@@
D.3 Example-based Infographic Chart Generation
In the user study, each participant was compensated with 30 USD for their participation. Fig. 28 illustrates the user study interface using a specific example. Figs. 29-30 present all 30 triplets of infographic charts: one reference, two infographic charts to be rated that are generated by GPT-Image-1 and our method, respectively.

### v5

@@ Appendix A Chart-rich Websites
Table 5: Chart-rich websites and their licenses. All websites in the table allow research use and the sharing of source URLs.
Name URL License
Statista statista.com CC BY-NC
Visual Capitalist visualcapitalist.com Customized
World Statistics world-statistics.org CC BY
Our World in Data ourworldindata.org CC BY
OECD oecd.org CC BY
Openverse openverse.org $C$
The Conversation theconversation.com CC BY-ND
Kaiser Family Foundation kff.org CC BY-NC-ND
Office for National Statistics ons.gov.uk OGL
Information is Beautiful informationisbeautiful.net Customized
Pew Research Center pewresearch.org Customized
MarketingCharts marketingcharts.com Customized
Chit Chart chitchart.com Customized
World Economic Forum weforum.org Customized
Wikimedia wikimediafoundation.org Customized
hikaku-sitatter hikaku-sitatter.com Customized
Eurostat ec.europa.eu Customized
European Parliament europarl.europa.eu Customized
Appendix B Pipeline for Data Table Extraction
For real infographic charts, we extract a per-chart data table using LVLMs in a multi-step verification pipeline, where humans are involved in resolving inconsistent tabular results. Specifically, we first run Gemini‑2.0‑Flash and GPT‑4o‑mini in parallel and retain tables where both models produce consistent outputs. Two tables are considered consistent if they have 1) the same number of data points, 2) identical column names and categorical values, and 3) numerical values differing by no more than 5%. The threshold of 5% is set by following the common practices (Methani et al., 2020; Masry et al., 2022; 2025). If the two tables are consistent but not exactly matched, we randomly select one for use. We also reviewed 200 randomly selected exact matches and found that all of them were correct. In contrast, if the two tables are inconsistent, we perform an additional verification step where GPT-4o also extracts a table. If GPT-4o’s output agrees with either of the models, we accept it. Otherwise, trained annotators from the service provider verify and correct the table. The success rates of this pipeline are:
  • 43.6% of charts are accepted directly when GPT‑4o‑mini and Gemini‑2.0‑Flash agree.
  • Among the remaining 56.4%, GPT‑4o resolves 76.2% of the inconsistencies, resulting in 86.6% of charts being automatically processed by LVLMs.
  • The remaining 13.4% are manually verified and corrected.
This pipeline greatly reduces the need for manual annotation while ensuring reliable outputs.
Appendix C Dataset Diversity
Geographic diversity We evaluated the geographic diversity of the whole dataset using Gemini-2.5-Flash to extract country and region references based on the UN geoscheme. Infographic charts with geographic information account for 59.4% of the total, with the most frequently referenced regions being North America (33.0%), Europe (23.1%), Asia (19.3%), and Africa (10.6%). While the dataset spans a broad range of regions, there is a moderate skew toward North America and Europe, reflecting the distribution of original sources (e.g., Visual Capitalist, Statista).
Domain diversity Likewise, we analyzed domain diversity in our dataset by prompting Gemini-2.5-Flash to assign each chart a topic based on the IPTC Media Topics taxonomy, a widely used hierarchical ontology for classifying news content. The result shows that the infographic charts in ChartGalaxy span 16 out of 17 top-level categories (excluding a sensitive category on religion conflict) and 99 out of 120 topics defined in the taxonomy. The most frequently occurring topics include public health (7.3%), business enterprise (6.0%), and environmental conservation (3.9%). These findings indicate that our dataset captures a wide and balanced distribution of real-world domains.
Linguistic diversity Following ChartQAPro (Masry et al., 2025), we measured linguistic diversity using the average pairwise cosine distance between Sentence-BERT embeddings extracted by the all-MiniLM-L6-v2 model (Reimers & Gurevych, 2019). Our dataset achieves a linguistic diversity score of 0.8754, higher than ChartQAPro (Masry et al., 2025) (0.8439) and Chartxiv (Wang et al., 2024) (0.7831). This result indicates that ChartGalaxy exhibits greater linguistic diversity than existing datasets. The real infographic charts predominantly use English, with additional content in languages such as Spanish, German, and French, reflecting the linguistic diversity of online data. All synthetic data in our dataset is in English, which can be easily translated into other languages upon request, offering flexibility to meet users’ diverse linguistic needs.
Representativeness of synthetic infographic charts To evaluate the distributional alignment of our synthetic infographic charts with real infographic charts, we conducted an additional quantitative analysis. Specifically, we extracted DreamSim (Fu et al., 2023) features (dimension = 1792) from both real and synthetic infographics, applied UMAP for projection with n_neighbors=100n\_neighbors=100, and measured grid-based coverage over a 20×20 spatial grid. We define coverage as the percentage of grid cells occupied by real infographic charts that are also covered by synthetic ones. The result shows that our synthetic charts cover 97.62% of the feature space occupied by real infographic charts, demonstrating that our synthetic infographics are highly representative of real infographics.
Appendix D Chart Types and Variations
We provide a full list of 75 chart types and 440 chart variations.
D.1 Chart Types
@@
We summarize chart types observed in the collected real infographic charts, drawing on two existing taxonomies: Data Viz Project and Datylon . The diversity of chart types ensures that our synthetic infographic charts can adapt to a wider range of scenarios and data representations, making them valuable for model training and evaluation. The full list of 75 chart types is in Figs. 5-20.
D.2 Chart Variations
For each chart type, we include multiple stylistic variations, designed along the following dimensions:
  • Element shapes, such as rounded bars and curved bars.
  • Icon placement, such as positioned above data elements and beside labels.
  • Rendering effects, such as hand-drawn style and 3D style.
  • Element alignment, such as center-aligned layout and edge-aligned layout.
  • Gridline and axis design, ranging from minimalist to detailed ticks and axes.
These variations were observed from real-world infographics and verified by design experts, who also added complementary styles to ensure coverage and diversity. This results in a variation space that reflects real-world design patterns and supports diverse chart generation. We provide illustrations of the variations under each chart type (Figs. 5-20).
Figure 5: 75 chart types and 440 chart variations (Part 1).
Figure 6: 75 chart types and 440 chart variations (Part 2).
Figure 7: 75 chart types and 440 chart variations (Part 3).
Figure 8: 75 chart types and 440 chart variations (Part 4).
Figure 9: 75 chart types and 440 chart variations (Part 5).
Figure 10: 75 chart types and 440 chart variations (Part 6).
Figure 11: 75 chart types and 440 chart variations (Part 7).
Figure 12: 75 chart types and 440 chart variations (Part 8).
Figure 13: 75 chart types and 440 chart variations (Part 9).
Figure 14: 75 chart types and 440 chart variations (Part 10).
Figure 15: 75 chart types and 440 chart variations (Part 11).
Figure 16: 75 chart types and 440 chart variations (Part 12).
Figure 17: 75 chart types and 440 chart variations (Part 13).
Figure 18: 75 chart types and 440 chart variations (Part 14).
Figure 19: 75 chart types and 440 chart variations (Part 15).
Figure 20: 75 chart types and 440 chart variations (Part 16).
Appendix E Chart Type Selection
@@
We determine chart types by analyzing the data attributes and their characteristics. First, we apply rule-based mapping (Table LABEL:tab:chart_mapping_rules) that identifies candidate chart types based on data attribute combinations. If multiple candidates remain, we instruct Gemini-2.0-Flash to select the best fit by evaluating the compatibility between the data and these chart types. The selection prompt shown below provides candidate chart types, their descriptions, and key data statistics to inform the decision.
@@
Appendix F Layout Templates
We summarize 68 layout templates from real infographic charts. The full list of these templates is shown in Figs. 21 and 22.
@@ Appendix G Examples of Synthetic Infographic Charts
Appendix H Extended Evaluation and Results
H.1 Instruction Dataset for Infographic Chart Understanding
@@
We fine-tune two LVLMs in our experiments. The first is InternVL3-8B (Zhu et al., 2025b), which combines the Qwen2.5-7B language model with the InternViT-300M-448px-V2.5 visual encoder. The second is Qwen2.5-VL-7B (Bai et al., 2025), which integrates the same Qwen2.5-7B language model with a Vision Transformer architecture optimized by Qwen. These models are fine-tuned on our instruction dataset for 2 epochs using a global batch size of 128. For parameter-efficient tuning, we utilize LoRA (Hu et al., 2022) with a rank of 8. This training process is conducted on 8 NVIDIA 4090D 48G GPUs, leveraging Deepspeed ZerO for parallelized training. For both InternVL3-8B and Qwen2.5-VL-7B, the learning rate is set to $5\times 10^{-5}$, and we employ a cosine learning rate scheduler with a warm-up ratio of $0.1$. The LoRA parameters, $\alpha$ and $r$, are both set to 32.
@@
Ablation Study on Real and Synthetic Data
To further understand the role of synthetic data in chart understanding, we conducted an ablation experiment on chart understanding. Specifically, we fine-tuned InternVL3-8B and Qwen2.5-VL-7B on the real and synthetic infographics separately, and compared their performance on three benchmarks: the ChartGalaxy evaluation set, ChartQAPro, and InfographicsVQA.
As shown in Table 8, the models fine-tuned on both real and synthetic infographics perform the best, with observable advantages over those fine-tuned solely on real or synthetic data. We further observe that the gains from synthetic data are particularly notable on our evaluation set. This improvement arises because synthetic infographics provide bounding-box annotations for charts, text, and images, which facilitates the generation of complex instructions, especially for visual-element-based reasoning.
Table 8: Comparison between real, synthetic, and combined data on three benchmarks.
Model ChartGalaxy evaluation set ChartQAPro InfographicsVQA
InternVL3-8B 53.20 38.15 76.19
+ Real 55.74 42.58 78.67
+ Synthetic 76.08 41.76 77.92
+ Both 80.07 44.13 79.99
Qwen2.5-VL-7B 56.50 37.97 78.59
+ Real 58.76 40.47 81.76
+ Synthetic 77.80 40.07 81.01
+ Both 80.35 41.56 83.03
Additional Analysis on a Smaller Model
We conducted additional experiments with a smaller model, Qwen2.5-VL-3B, and fine-tuned it on the ChartGalaxy QA dataset (Sec. 4.1). The evaluation was performed on three benchmarks: the ChartGalaxy evaluation set, ChartQAPro, and InfograpicVQA.
Table 9: Evaluation of Qwen2.5-VL-3B with and without fine-tuning on ChartGalaxy QA dataset.
Model ChartGalaxy evaluation set ChartQAPro InfograpicVQA
Qwen2.5-VL-3B 52.42 30.89 74.39
+ ChartGalaxy 71.01 37.08 78.43
The results show consistent improvements across all benchmarks after fine-tuning. Notably, the model achieves a +18.59 gain on the ChartGalaxy evaluation set and +6.19 on ChartQAPro, demonstrating that smaller models can also effectively benefit from training on ChartGalaxy. These findings suggest that our dataset is broadly applicable and valuable even for resource-constrained VLMs.
H.2 Benchmarking Infographic Chart Code Generation
@@
Table 11 lists the 17 LVLMs along with their API names used in this experiment. Since our task involves generating executable code with relatively long outputs, we conducted a small-scale pilot study to assess the basic code generation capability of each model. Based on this, we excluded models that consistently failed to produce meaningful or complete outputs under our task setting—for example, Phi-4—due to their limited capacity or inability to handle long sequences. For illustration, we present the results of two smaller models (Qwen2.5-VL-7B, Phi-4-6B) on our benchmark in Table 10. We use greedy decoding (temperature $\tau=0$) across all models to ensure deterministic outputs. To maximize the chance of obtaining complete and executable code, we configure each model to generate as many tokens as possible, setting the maximum generation length to $min(16384,A)$, where $A$ denotes the model’s maximum generation limit. This helps mitigate the risk of incomplete outputs, which result in non-executable code.
Table 10: Evaluation results of Qwen2.5-VL-7B and Phi-4-6B.
Model Exec. Rate Low-Level High-Level Overall
Qwen2.5-VL-7B 43.80 13.85 9.25 11.55
Phi-4-6B 1.60 0.29 0.07 0.18
@@
Gemini-2.5-Pro (Google, 2025b) Proprietary gemini-2.5-pro-preview-05-06
Gemini-2.5-Flash (Google, 2025a) Proprietary gemini-2.5-flash-preview-04-17
Claude-3.7-Sonnet (Anthropic, 2025) Proprietary claude-3-7-sonnet-20250219
GPT-4.1 (OpenAI, 2025a) Proprietary gpt-4.1
GPT-4.1-mini (OpenAI, 2025a) Proprietary gpt-4.1-mini
GPT-4.1-nano (OpenAI, 2025a) Proprietary gpt-4.1-nano
OpenAI-o4-mini (OpenAI, 2025d) Proprietary o4-mini
OpenAI-o3 (OpenAI, 2025d) Proprietary o3
OpenAI-o1 (OpenAI, 2025c) Proprietary o1
GPT-4o (OpenAI, 2025b) Proprietary gpt-4o-2024-11-20
Doubao-1.5-Vision-Pro (ByteDance, 2025) Proprietary Doubao-1.5-vision-pro-32k
Moonshot-v1-Vision (Moonshot AI, 2025) Proprietary moonshot-v1-32k-vision-preview
Llama-4-Maverick-17B (Meta, 2025) Open-Source chutesai/Llama-4-Maverick-17B-128E-Instruct-FP8
Llama-4-Scout-17B (Meta, 2025) Open-Source chutesai/Llama-4-Scout-17B-16E-Instruct
Qwen2.5-VL-72B (Bai et al., 2025) Open-Source Qwen/Qwen2.5-VL-72B-Instruct
Qwen2.5-VL-32B (Bai et al., 2025) Open-Source Qwen/Qwen2.5-VL-32B-Instruct
InternVL3-78B (Zhu et al., 2025b) Open-Source internvl3-78b
@@
For the high-level score, we employ GPT-4o (OpenAI, 2025b) to assess the visual similarity between the PNG image rendered by the generated code and the ground-truth one. We instruct GPT-4o to evaluate the similarity along six dimensions: data element, layout, text, image, color, and validity. The model outputs a score for each of the six dimensions, which are then summed to produce a total score ranging from 0 to 100. The detailed prompt for this evaluation is provided in Supp. I.2.
The low-level score evaluates the fine-grained similarity between SVG elements of the rendered chart and the corresponding ground-truth chart. This evaluation is conducted through three steps: 1) decomposing both charts into SVG elements, 2) matching elements between the two charts, and 3) computing similarity metrics based on the matching results (Si et al., 2025; Chen et al., 2024). Algorithm 1 presents the pseudo-code for the matching procedure. The full implementation details are available in our publicly accessible code repository [^2] https://github.com/ChartGalaxy/ChartGalaxy.
@@
1:gt_leafsgt\_leafs: Ground truth SVG elements.
2:pr_leafspr\_leafs: Predicted SVG elements.
3:gt_matchedgt\_matched: Array for ground truth matches (init with -1).
4:pr_matchedpr\_matched: Array for prediction matches (init with -1).
@@
[…]
@@
[…]
@@
15:  if $CostMatrix[i][j]\leq 1$ AND gt_matched[i]=1gt\_matched[i]=-1 AND pr_matched[j]=1pr\_matched[j]=-1 then
@@
The text and image metrics evaluate the similarity of generated text and image elements, respectively, by averaging the similarity scores over all matched pairs of text and image elements. Unmatched text and image elements in ground-truth charts are assigned a similarity score of 0 to penalize generation failures. For text elements, similarity is computed using the character-level Sørensen-Dice coefficient, defined as twice the number of overlapping characters divided by the total number of characters in the two strings (Si et al., 2025). For image elements, similarity is measured using the CLIP embedding-based similarity (Radford et al., 2021).
The color, position, and size metrics assess visual consistency across matched elements with respect to their respective attributes. The color metric employs the CIEDE2000 formula (Luo et al., 2001) to measure the perceptual difference between the colors of matched elements. The position and size metrics are defined as follows:
@@
Gemini-2.5-Pro 6,662.45 1,674.01 85.21
GPT-4.1 5,237.78 1,675.36 80.00
Claude-3.7-Sonnet 5,870.50 1,552.48 79.91
GPT-4.1-mini 7,656.71 2,488.46 79.69
OpenAI-o4-mini 2,886.75 735.19 75.97
Gemini-2.5-Flash 6,508.22 1,934.79 75.55
OpenAI-o1 2,662.11 1,018.11 74.69
OpenAI-o3 3,114.74 999.66 74.22
GPT-4o 2,962.72 613.42 65.67
GPT-4.1-nano 2,387.05 1,306.72 60.06
Doubao-1.5-Vision-Pro 3,878.70 1,347.10 47.11
Moonshot-v1-Vision 3,087.17 791.16 44.39
@@
Llama-4-Maverick-17B 3,460.76 856.62 61.29
Qwen2.5-VL-72B 3,512.20 1,236.29 57.09
InternVL3-78B 3,376.52 758.71 55.07
Llama-4-Scout-17B 3,247.25 826.56 51.91
Qwen2.5-VL-32B 4,960.82 1,264.43 46.48
@@
Prompting methods. In addition to the direct prompting method, we evaluate three alternative prompting strategies following prior work (Yang et al., 2025; Si et al., 2025):
@@
  • SelfReflection, where the model is provided the ground-truth chart, previously generated code from direct prompting, and the rendered chart, and is instructed to revise the generated code.
@@
Llama-4-Maverick-17B Direct 100.00 78.24 58.51 59.53 66.35 74.00 47.74 64.06 58.42 61.24
@@
H.3 Example-based Infographic Chart Generation
The study was approved by the Institutional Review Board of the first author’s university. In the user study, each participant was compensated with 30 USD for their participation. The study did not involve exposure to emotionally charged, political, or misleading content. Participants were shown infographic charts on neutral topics (e.g., bird population growth, energy sources) and asked to evaluate their quality. No sensitive or personally identifiable data was collected during the study. Participants were fully informed of their rights and were free to withdraw at any time. Fig. 28 illustrates the user study interface using a specific example. Figs. 29-30 present all 30 triplets of infographic charts: one reference, two infographic charts to be rated that are generated by GPT-Image-1 and our method, respectively.